# SegNet4D: Efficient Instance-Aware 4D LiDAR Semantic Segmentation for Driving Scenarios

### v2

@@ Abstract
4D LiDAR semantic segmentation, also referred to as multi-scan semantic segmentation, plays a crucial role in enhancing the environmental understanding capabilities of autonomous robots. It classifies the semantic category of each LiDAR point and detects whether it is dynamic, a critical ability for tasks like obstacle avoidance and autonomous navigation. Existing approaches often rely on computationally heavy 4D convolutions or recursive networks, which result in poor real-time performance, making them unsuitable for online robotics and autonomous driving applications. In this paper, we introduce SegNet4D, a novel real-time 4D semantic segmentation network offering both efficiency and strong semantic understanding. SegNet4D addresses 4D segmentation as two tasks: single-scan semantic segmentation and moving object segmentation, each tackled by a separate network head. Both results are combined in a motion-semantic fusion module to achieve comprehensive 4D segmentation. Additionally, instance information is extracted from the current scan and exploited for instance-wise segmentation consistency. Our approach surpasses state-of-the-art in both multi-scan semantic segmentation and moving object segmentation while offering greater efficiency, enabling real-time operation. Besides, its effectiveness and efficiency have also been validated on a real-world robotic platform. Our code will be released at https://github.com/nubot-nudt/SegNet4D.
@@
4D semantic segmentation, deep learning, moving object segmentation.I Introduction
@@
Light Detection and Ranging (LiDAR) sensors have become important for autonomous vehicles and robots to achieve autonomy due to their resilience to illumination changes, wide field of views, and accurate range measurements [1, 2]. As a result, LiDAR-based semantic perception has garnered extensive research interest in recent years. Semantic segmentation using 3D LiDAR data aims to assign a specific semantic class label to each point in the acquired point clouds, serving as a fundamental task in 3D perception and scene understanding. It plays a critical role in enhancing downstream tasks such as point cloud registration [3] and simultaneous localization and mapping (SLAM) [4].
@@
In practical applications, however, single-scan 3D LiDAR semantic segmentation methods [1, 5, 6] are limited because they lack motion information, which is critical for obstacle avoidance and path planning in autonomous robotic and driving applications. To address this, 4D semantic understanding is necessary, as it categorizes each point while identifying its dynamic attributes [7, 8, 9]. This task typically relies on sequential LiDAR scans to capture motion information. In this regard, some methods directly stack historical LiDAR scans into a single pointcloud and feed it into a single-scan-based network for achieving multi-scan segmentation [1, 5, 10], which leads to suboptimal performance due to the lack of temporal information association. Other approaches use 4D convolutional neural networks [11] or recursive networks [8, 12] to extract motion features from input sequential LiDAR scans. However, these approaches impose heavy computational burdens and make real-time operations challenging. Moreover, existing methods commonly treat multi-scan semantic segmentation in an end-to-end fashion, directly predicting semantic labels for all categories, including those moving and static classes. This may lead to limited performance in recognizing dynamic objects because moving points are often fewer than static background points in most datasets.
@@
To address these challenges, building upon our previous work InsMOS [13], we propose an extended framework for 4D semantic segmentation. Our key idea is to tackle 4D semantic segmentation by dividing it into two subtasks: single-scan semantic segmentation (SSS) and moving object segmentation (MOS). We argue that semantic classes and motion characteristics represent distinct aspects of objects. By addressing them separately and then fusing their outputs, our method achieves a more comprehensive understanding and superior performance compared to end-to-end approaches. To enhance efficiency, our framework converts the sequential LiDAR scans into Bird’s Eye View (BEV) images and extracts motion features by calculating BEV residuals, significantly reducing computational cost compared to existing fashion using 4D convolutions. Furthermore, we find instance information [14, 13] is crucial for the LiDAR segmentation tasks to prevent over-segmentation. However, existing methods struggle to integrate this information effectively. To address this, we incorporate instance consistency from the current scan into the prediction pipeline at both feature and point levels, enabling instance-aware segmentation. Finally, we design a novel module to fuse point-wise semantic predictions and motion states to achieve accurate 4D semantic segmentation online. We extensively evaluate SegNet4D on mainstream datasets, comparing its performance with state-of-the-art (SOTA) methods. The results confirm that our approach surpasses SOTA in both multi-scan semantic segmentation and MOS while supporting real-time operation. Additionally, we validate its practical utility on a real robot platform, demonstrating both its effectiveness and efficiency for autonomous driving applications.
@@
We propose a novel 4D semantic segmentation framework that decomposes the task into two subtasks: Single-Scan Semantic Segmentation (SSS) and Moving Object Segmentation (MOS). The results are then combined using a motion-semantic fusion module, achieving SOTA performance in both 4D semantic segmentation and MOS.
@@
We propose an efficient 4D semantic segmentation network that is significantly faster than existing methods and supports real-time operation. Furthermore, we successfully integrate the network into a real robot, demonstrating its practical utility for self-driving applications.
This article is an extension of our previous conference paper, i.e., InsMOS [13], which proposed to add instance information for MOS. Compared to [13], this article extends it in four critical aspects: 1) a fully 4D semantic segmentation method that predicts both the motion states and semantic categories. 2) a novel module for integrating point-wise motion states and semantic category predictions. 3) a more efficient fashion for encoding motion features, enabling this method to operate in real-time. 4) a more extensive experimental evaluation on multiple datasets and a real-world robot platform, demonstrating this method’s superior performance and practical utility.
@@ II Related Work
MOS refers to accurately identifying dynamic regions or objects from input data. The task has attracted extensive research in the field of computer vision [15, 16]. In recent years, some research has also been paid attention to LiDAR-based MOS. We categorize these methods into projection-based and non-projection-based.
The projection-based method typically converts sequential point clouds into image representations, such as range images [17, 2, 18, 19] or BEV images [20, 21], and then compute the residuals of sequential images to extract motion information in the scene. Chen et al. [17] first released the MOS datasets and benchmark based on the SemanticKITTI dataset [10], and proposed a deep neural network to learn motion cues from sequential residual range images for online MOS. To extract motion features more comprehensively, Sun et al. [18] proposed a dual-branch network architecture that separately encodes the appearance features and temporal motion features, and then fuses them with a multi-scale motion-guided attention module. Subsequently, Cheng et al. [19] utilized a similar dual-branch architecture and proposed a distribution-based data augmentation method to improve the network robustness. Kim et al. [2] also designed a similar multi-branch network to improve MOS accuracy. Unlike range images, BEV images offer a top-down perspective, allowing a more intuitive observation of object movement. Based on this, Mohapatra et al. [20] designed a lightweight network architecture for extracting motion information from BEV residual images and successfully achieved real-time operation in an embedded platform. However, this method can only perform pixel-wise segmentation in the BEV space and has relatively low accuracy. Recently, Zhou et al. [21] projected 3D point clouds into the polar coordinate BEV space and designed a dual-branch network structure similar to [18] to enhance the MOS performance.
@@ III-A Motion Feature Encoding Module
4D semantic segmentation not only predicts the semantic category for each point but also identifies its motion state. Therefore, it is necessary to extract motion features from the sequential point clouds. Existing methods mainly utilize 4D convolution [22, 23] to obtain motion cues, which poses a significant computational workload. To improve the real-time performance, we adopt the BEV image representation for fast motion feature encoding, as shown in Fig. 2, which avoids computational expense processing on extensive unstructured point clouds data. The encoding process is divided into three steps as follows.
Point Cloud Alignment. To compensate for the ego-motion of LiDAR, spatial alignment is conducted initially on the input sequential point clouds. Specifically, given the current point cloud consisting of points represented in homogeneous coordinates as , along with the past consecutive point clouds with their transformations relative to , we transform the past consecutive point clouds into the current viewpoint by
@@ IV Experimental Evaluation
In this section, we conduct a series of experiments on SemanticKITTI [10] and nuScenes [37] datasets to demonstrate SegNet4D’s capabilities on multi-scan semantic segmentation (Sec. IV-A) and MOS (Sec. IV-B) and compare the performance with SOTA. Subsequently, we integrate our methodology into a real robot platform (Sec. IV-C), supporting online real-time operation. Besides, we perform ablation experiments on our network to evaluate the effectiveness of the framework and the proposed modules (Sec. IV-D). These experimental results will substantiate our claims regarding the contributions.
Datasets-SemanticKITTI [10]. SemanticKITTI provides semantic labels for each individual LiDAR scan in the KITTI [38] odometry dataset, which comprises 22 sequences collected with a Velodyne HDL-64E LiDAR. Following previous work [1, 24], we use the standard data split where sequences 00 to 10 are used for training (with sequence 08 for validation), and sequence 11 to 21 for testing. The dataset contains multiple semantic category labels. In our model, the training is supervised with 26 semantic categories, including 6 moving classes, 19 static classes, and one class. For the MOS task, All the semantic categories are reorganized into two classes: moving and static [17, 22]. Due to the imbalanced distribution of moving objects between the training and test set, an additional dataset, KITTI-Road, is introduced in [18] to mitigate the impact. We follow the experimental setups described in [18] to test the MOS performance. Datasets-nuScenes [37]. It consists of 1000 driving scenes collected with a 32-beam LiDAR sensor. For point cloud segmentation tasks, it is mainly used for evaluating SSS with 16 static semantic classes. To achieve multi-scan segmentation, we utilize the annotated bounding box motion attributes to generate 8 new moving categories, including , , , , , , , and . We will release the multi-scan semantic segmentation dataset of nuScenes for the convenience of the community. Similar to the SemanticKITTI, we divide all semantic categories into two classes (moving and static) for evaluating MOS.
@@ IV-B Evaluation for Moving Object Segmentation
IV-C Operation on Real Robots
To demonstrate the real-world utility of SegNet4D, we further simplify our model and integrate it into a robotic system. As shown in Fig. 7, our robot platform is equipped with a RoboSense Helios-32 LiDAR with 10 Hz sampling frequency, an Xsens MTI-300 IMU, a CMOS camera, and an industrial computer with Intel Core i7-1165G7, 16G RAM, and NVIDIA RTX 2060 GPU.
Our method achieves real-time operation at 15.7 Hz rate on the self-developed platform, faster than the typical rotation LiDAR sensor frame-rate of 10 Hz. Its parameter count is merely 35.7 M, indicating small memory consumption suitable for limited robot onboard resources. We also utilize this platform to collect a dataset from the campus scene and qualitatively compare the semantic segmentation results with Cylinder3D [1] and Mars3D [9]. All methods are only trained on the nuScenes dataset and directly applied for inference on our dataset. As illustrated in Fig. 8, we can see that our SegNet4D gets the best segmentation results, demonstrating its practical utility for the robot platform. Additionally, we manually annotate a sequence (900 frames) with moving labels to quantitatively evaluate the MOS generalization performance. From the results shown in Tab. IV, our method still possesses the best generalization ability.
The above-mentioned experimental results show that SegNet4D can operate in real-time on the real robot platform while exhibiting superior performance in 4D semantic segmentation and MOS. This highlights its practical utility in enhancing robots’ environmental perception capabilities.
@@ V Conclusion
In this paper, we present a novel 4D semantic segmentation method to predict both point-wise moving labels and semantic labels for LiDAR data, and operate in real-time. The framework decomposes the complex 4D semantic segmentation task into MOS and SSS tasks, finally merging their predictions to achieve more accurate 4D semantic segmentation. We adopt a projection-based approach to quickly obtain motion features, which significantly reduces the computational complexity compared to 4D convolutions. To achieve instance-aware segmentation, we concatenate the motion features with the spatial features of the current scan, feeding them into the network for instance detection, and subsequently inject the instance features into the prediction pipeline. In addition, we design a motion-semantic fusion module to integrate the point-wise motion states and static semantic predictions, enabling motion-guided 4D semantic segmentation. Extensive experiments on multiple datasets and a real-world robot platform demonstrate that our approach is effective and efficient.

### v3

# SegNet4D: Efficient Instance-Aware 4D Semantic Segmentation for LiDAR Point Cloud

@@ Abstract
4D LiDAR semantic segmentation, also referred to as multi-scan semantic segmentation, plays a crucial role in enhancing the environmental understanding capabilities of autonomous vehicles or robots. It classifies the semantic category of each LiDAR measurement point and detects whether it is dynamic, a critical ability for tasks like obstacle avoidance and autonomous navigation. Existing approaches often rely on computationally heavy 4D convolutions or recursive networks, which result in poor real-time performance, making them unsuitable for online robotics and autonomous driving applications. In this paper, we introduce SegNet4D, a novel real-time 4D semantic segmentation network offering both efficiency and strong semantic understanding. SegNet4D addresses 4D segmentation as two tasks: single-scan semantic segmentation and moving object segmentation, each tackled by a separate network head. Both results are combined in a motion-semantic fusion module to achieve comprehensive 4D segmentation. Additionally, instance information is extracted from the current scan and exploited for instance-wise segmentation consistency. Our approach surpasses state-of-the-art in both multi-scan semantic segmentation and moving object segmentation while offering greater efficiency, enabling real-time operation. Besides, its effectiveness and efficiency have also been validated on a real-world unmanned ground platform. Our code will be released at https://github.com/nubot-nudt/SegNet4D.
@@
4D semantic segmentation, deep learning, moving object segmentation, LiDAR point cloud.I Introduction
@@
Light Detection and Ranging (LiDAR), employing time of flight (ToF) measurement technology [1], enables accurately capturing environmental geometry details. Known for its resilience to illumination changes, wide field of views, and accurate measurement, LiDAR have become an important instrument for autonomous vehicles and robots. As a result, LiDAR-based semantic perception has garnered extensive research interest in recent years. Semantic segmentation using 3D LiDAR data aims to assign a specific semantic class label to each point in the acquired environmental measurements, serving as a fundamental task in 3D perception and scene understanding. It plays a critical role in enhancing downstream tasks such as point cloud registration [2] and simultaneous localization and mapping (SLAM) [3].
@@
In practical applications, however, single-scan 3D LiDAR semantic segmentation methods [4, 5, 6, 7] are limited because they lack motion information, which is critical for obstacle avoidance and path planning in autonomous robotic and driving applications. To address this, 4D semantic understanding is necessary, as it categorizes each LiDAR measurement while identifying its dynamic attributes [8, 9, 10]. This task typically relies on sequential LiDAR scans to capture motion information. In this regard, some methods directly stack historical LiDAR scans into a single pointcloud and feed it into a single-scan-based network for achieving multi-scan segmentation [4, 6, 11], which leads to suboptimal performance due to the lack of temporal information association. Other approaches use 4D convolutional neural networks [12] or recursive networks [9, 13] to extract motion features from input sequential LiDAR scans. However, these approaches impose heavy computational burdens and make real-time operations challenging. Moreover, existing methods commonly treat multi-scan semantic segmentation in an end-to-end fashion, directly predicting semantic labels for all categories, including those moving and static classes. This may lead to limited performance in recognizing dynamic objects because moving points are often fewer than static background points in most datasets.
@@
To address these challenges, building upon our previous work InsMOS [14], we propose an extended framework for 4D semantic segmentation. Our key idea is to tackle 4D semantic segmentation by dividing it into two subtasks: single-scan semantic segmentation (SSS) and moving object segmentation (MOS). We argue that semantic classes and motion characteristics represent distinct aspects of objects. By addressing them separately and then fusing their outputs, our method achieves a more comprehensive understanding and superior performance compared to end-to-end approaches. To enhance efficiency, our framework converts the sequential LiDAR scans into Bird’s Eye View (BEV) images and extracts motion features by calculating BEV residuals, significantly reducing computational cost compared to existing fashion using 4D convolutions. Furthermore, we find instance information [15, 14] is crucial for the LiDAR segmentation tasks to prevent over-segmentation. However, existing methods struggle to integrate this information effectively. To address this, we incorporate instance consistency from the current scan into the prediction pipeline at both feature and point levels, enabling instance-aware segmentation. Finally, we design a novel module to fuse point-wise semantic predictions and motion states to achieve accurate 4D semantic segmentation online. We extensively evaluate SegNet4D on mainstream datasets, comparing its performance with state-of-the-art (SOTA) methods. The results confirm that our approach surpasses SOTA in both multi-scan semantic segmentation and MOS while supporting real-time operation. Additionally, we validate its practical utility on a real unmanned ground platform, demonstrating both its effectiveness and efficiency for autonomous applications.
@@
We propose a novel 4D semantic segmentation framework for LiDAR measurements, which decomposes the task into two subtasks: Single-Scan Semantic Segmentation (SSS) and Moving Object Segmentation (MOS). The results are then combined using a motion-semantic fusion module, achieving SOTA performance in both 4D semantic segmentation and MOS.
@@
We propose an efficient 4D semantic segmentation network that is significantly faster than existing methods and supports real-time operation. Furthermore, we successfully integrate the network into a real platform, demonstrating its practical utility for autonomous applications.
This article is an extension of our previous conference paper, i.e., InsMOS [14], which proposed to add instance information for MOS. Compared to [14], this article extends it in four critical aspects: 1) a fully 4D semantic segmentation method that predicts both the motion states and semantic categories. 2) a novel module for integrating point-wise motion states and semantic category predictions. 3) a more efficient fashion for encoding motion features, enabling this method to operate in real-time. 4) a more extensive experimental evaluation on multiple datasets and a real-world unmanned ground platform, demonstrating this method’s superior performance and practical utility.
@@ II Related Work
LiDAR-based MOS refers to accurately identifying motion state for each measurement point. Existing methods can be divided into two groups, projection-based and non-projection-based fashion.
The projection-based method typically converts sequential point clouds into image representations, such as range images [16, 17, 18, 19] or BEV images [20, 21], and then compute the residuals of sequential images to extract motion information in the scene. Chen et al. [16] first released the MOS datasets and benchmark based on the SemanticKITTI dataset [11], and proposed a deep neural network to learn motion cues from sequential residual range images for online MOS. Subsequently, some work [18, 17, 19] attempt to enhance the network through encodes the appearance features and temporal motion features separately, and then fuses them with a multi-scale motion-guided attention module. Unlike range images, BEV images offer a top-down perspective, allowing a more intuitive observation of object movement. Based on this, Mohapatra et al. [20] designed a lightweight network architecture for extracting motion information from BEV residual images and successfully achieved real-time operation in an embedded platform. However, this method can only perform pixel-wise segmentation in the BEV space and has relatively low accuracy. Recently, Zhou et al. [21] projected 3D point clouds into the polar coordinate BEV space and designed a dual-branch network structure similar to [18] to enhance the MOS performance.
@@ III-A Motion Feature Encoding Module
4D semantic segmentation not only predicts the semantic category for each point measured by LiDAR but also identifies its motion state. Therefore, it is necessary to extract motion features from the sequential point clouds. Existing methods mainly utilize 4D convolution [22, 23] to obtain motion cues, which poses a significant computational workload. To improve the real-time performance, we adopt the BEV image representation for fast motion feature encoding, as shown in Fig. 2, which avoids computational expense processing on extensive unstructured point clouds data. The encoding process is divided into three steps as follows.
Point Cloud Alignment. To compensate for the ego-motion of LiDAR, spatial alignment is conducted initially on the input sequential point clouds. Specifically, given the current point cloud consisting of points represented in homogeneous coordinates as , along with the past consecutive point clouds with their relative transformations , we transform the past consecutive point clouds into the current viewpoint by
@@ IV Experimental Evaluation
In this section, we conduct a series of experiments on SemanticKITTI [11] and nuScenes [37] datasets to demonstrate SegNet4D’s capabilities on multi-scan semantic segmentation (Sec. IV-A) and MOS (Sec. IV-B) and compare the performance with SOTA. Subsequently, we integrate the method into a real unmanned ground platform (Sec. IV-C), supporting online real-time operation. Besides, we perform ablation experiments to evaluate the effectiveness of the framework and the proposed modules (Sec. IV-D). These experimental results will substantiate our claims regarding the contributions.
Datasets-SemanticKITTI [11]. SemanticKITTI provides semantic labels for each individual LiDAR scan in the KITTI [38] odometry dataset, which comprises 22 sequences collected with a Velodyne HDL-64E LiDAR. Following previous work [4, 24], we use the standard data split where sequences 00 to 10 are used for training (with sequence 08 for validation), and sequence 11 to 21 for testing. The dataset contains multiple semantic category labels. In our model, the training is supervised with 26 semantic categories, including 6 moving classes, 19 static classes, and one class. For the MOS task, All the semantic categories are reorganized into two classes: moving and static [16, 22]. Due to the imbalanced distribution of moving objects between the training and test set, an additional dataset, KITTI-Road, is introduced in [18] to mitigate the impact. We follow the experimental setups described in [18] to test the MOS performance.
Datasets-nuScenes [37]. It consists of 1000 driving scenes collected with a 32-beam LiDAR sensor. For point cloud segmentation tasks, it is mainly used for evaluating SSS with 16 static semantic classes. To achieve multi-scan segmentation, we utilize the annotated bounding box motion attributes to generate 8 new moving categories, including , , , , , , , and . We will release the multi-scan semantic segmentation dataset of nuScenes for the convenience of the community. Similar to the SemanticKITTI, we divide all semantic categories into two classes (moving and static) for evaluating MOS.
@@ IV-B Evaluation for Moving Object Segmentation
IV-C Operation on Real-World Platform
To demonstrate the real-world utility of SegNet4D, we further simplify our model and integrate it into a unmanned ground system. As shown in Fig. 7, our platform is equipped with a RoboSense Helios-32 LiDAR with 10 Hz sampling frequency, an Xsens MTI-300 IMU, a CMOS camera, and an industrial computer with Intel Core i7-1165G7, 16G RAM, and NVIDIA RTX 2060 GPU.
Our method achieves real-time operation at 15.7 Hz rate on the self-developed platform, faster than the typical rotation LiDAR sensor frame-rate of 10 Hz. Its parameter count is merely 35.7 M, indicating small memory consumption suitable for limited onboard resources. We also utilize this platform to collect a dataset from the campus scene and qualitatively compare the semantic segmentation results with Cylinder3D [4] and Mars3D [10]. All methods are only trained on the nuScenes dataset and directly applied for inference on our dataset. As illustrated in Fig. 8, we can see that our SegNet4D gets the best segmentation results, demonstrating its practical utility for the real platform. Additionally, we manually annotate a sequence (900 frames) with moving labels to quantitatively evaluate the MOS generalization performance. From the results shown in Tab. IV, our method still possesses the best generalization ability.
The above-mentioned experimental results show that SegNet4D can operate in real-time on the real unmanned ground platform while exhibiting superior performance in 4D semantic segmentation and MOS. This highlights its practical utility in enhancing robots’ environmental perception capabilities.
@@ V Conclusion
In this paper, we present a novel 4D semantic segmentation method to predict both point-wise moving labels and semantic labels for LiDAR measurements, and operate in real-time. The framework decomposes the complex 4D semantic segmentation task into MOS and SSS tasks, finally merging their predictions to achieve more accurate 4D semantic segmentation. We adopt a projection-based approach to quickly obtain motion features, which significantly reduces the computational complexity compared to 4D convolutions. To achieve instance-aware segmentation, we concatenate the motion features with the spatial features of the current scan, feeding them into the network for instance detection, and subsequently inject the instance features into the prediction pipeline. In addition, we design a motion-semantic fusion module to integrate the point-wise motion states and static semantic predictions, enabling motion-guided 4D semantic segmentation. Extensive experiments on multiple datasets and a real-world unmanned ground platform demonstrate that our approach is effective and efficient.